OpenStudy (anonymous):

Type of snack Number sold Energy bars 194 Baked chips 200 Organic treats 235 Whole grain crackers 78 The club sold the snacks for $3 each. The club gets half of this amount as profit. The club also gets a bonus of $100 for any snack when it sells at least 200 of that snack. Last year’s data on sales are shown in the table. (b) Write an inequality to show how many snacks of one type would need to be sold to earn a profit of at least $500. Support your inequality with an explanation.

OpenStudy (anonymous):

hi can you help?

OpenStudy (loser66):

No, I don't get the question.

OpenStudy (anonymous):

oh

OpenStudy (loser66):

$3/each, profit is half of this amount + $100 bonus if selling more than 200 packs. So that, for baked chips only, the shop sells 200 pack, get bonus $100 and the profit is $150. Total is $250 profit for baked chips only. for Organic treat, it gets $100 for bonus, profit is 235*3 /2 = $352.50 Total is 452.50 for Organic treat only. Just 2 of them as above, the profit exceed$500. But in general, the profit for each of them is \(P(x) =\dfrac{3x}{2}+100~such~that ~ 200\leq x\)

OpenStudy (loser66):

Combine all, we have \(P(x) =\begin{cases}\dfrac{3x}{2} ~~~if ~~x <200\\\dfrac{3x}{2}+100~~if~~x \geq 200\end{cases}\)